OpenStudy (anonymous):

How can you determine algebraically that the points A(1,2), B(2,4), C(3,6) are on the same line?

OpenStudy (amistre64):

define line?

OpenStudy (anonymous):

Graph x and y axis

OpenStudy (amistre64):

if they have the same difference between them; they are on the same slope

OpenStudy (amistre64):

( 1, 2) +1 +2 ------ ( 2 , 4) +1 +2 ------ ( 3 , 6)

OpenStudy (amistre64):

if : \[\frac{y_1-y_2}{x_1-x_2}=\frac{y_2-y_3}{x_2-x_3}\] then they are on the same line

OpenStudy (anonymous):

you can also visualize. from (1,2) to (2,4) is over one unit, up two. (slope is 2) then from (2,4) to (3,6) is also over one up two. so yes, on the same line

OpenStudy (amistre64):

algebra aint visual lol

OpenStudy (anonymous):

it is when you are drawing lines yes?

OpenStudy (amistre64):

lines are analytic geometry; now do we 'prove' it algebraically?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

when m=m and they share a common point; its proofed

OpenStudy (anonymous):

You two stop fighting!! and @amistre: Having the same slope does not necessarily means that they are on the same line. Right?!

OpenStudy (anonymous):

yes of course you are right (master) but some thinks are obviously so. i know for a fact that if you had two points (1,2) and (2,4) you would find the slope with your eyeballs, not by subtracting and dividing

OpenStudy (anonymous):

mean*

OpenStudy (amistre64):

it means they are on the same line f they share a common point right?

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

yes of course. over 1 up 2, over 1 up 2, over 1 up 2 ...

OpenStudy (anonymous):

wish i knew what an inverse transform was. guess i should look it up

OpenStudy (anonymous):

In this example, it's so clear that \(y=2x\). So, yes they are on the same line. But if we have much more complicated points, we need to find an equation for the line using two points and then see if the third point lie on the same line by substituting in the equation.

OpenStudy (amistre64):

'How can you determine algebraically that .....' ; draw a picture lol

OpenStudy (anonymous):

@satellite: What inverse transforms are you talking about?